OpenStudy (anonymous):

what is this 75x3−65x2+10x factored completely

OpenStudy (anonymous):

75x^3-65x^2+10x

OpenStudy (mathlegend):

First think what can all these numbers be divided by?

OpenStudy (anonymous):

5

OpenStudy (mathlegend):

Good and all the numbers have the variable 'x' next to them right? So lets say 5x

OpenStudy (mathlegend):

So lets factor that out

OpenStudy (mathlegend):

75/5 = ? -65/5 = ? 10/5 = ?

OpenStudy (anonymous):

its 5x(15x^2-13x+2)

OpenStudy (mathlegend):

Good

OpenStudy (mathlegend):

Now what I like to use from here is what is known as the AC method

OpenStudy (mathlegend):

A= 15 B= -13 C= 2 A*C= ? (When multiplying 15 and 2) +B= ? (Rewrite 13... but think when using those two numbers above when does it equal -13)

OpenStudy (mathlegend):

Do you understand that part?

OpenStudy (anonymous):

kind of

OpenStudy (mathlegend):

Okay well in the parentheses we had 15x^2-13x+2 right? Now in textbooks this is written with letters like this.."ax^2+bx+c"

OpenStudy (mathlegend):

So you see where I pulled A, B, and C from. 15x^2-13x+2 ax^2+bx+c

OpenStudy (anonymous):

so is 15x^2-13x+2 the answer? :0

OpenStudy (mathlegend):

No

OpenStudy (anonymous):

what it, i can work backwords from the answer at this point and i think understand better

OpenStudy (mathlegend):

A= 15 B= -13 C= 2 A*C= 30 +B= -13 The two numbers that add to -13 but when multiplied that equal 30 are what?

OpenStudy (anonymous):

-10 -3

OpenStudy (mathlegend):

Okay so when we rewrite the equation what do you get? 15x^2-3x-10x+2 right? Notice I just split up the -13

OpenStudy (anonymous):

yess i get that

OpenStudy (mathlegend):

15x^2-3x-10x+2 So now find the common factors. Take the two first numbers and factor those. Then the last two. 15x^2-3x (Can you factor that?) -10x+2 (Can you factor that?)

OpenStudy (mathlegend):

Okay for 15x^2-3x the greatest number that goes into both of them is what?

OpenStudy (anonymous):

3

OpenStudy (mathlegend):

Good, but lets say 3x since both numbers have at least one 'x'. 15/3 = ? -3/3 = ?

OpenStudy (anonymous):

what is the answer to it? i am really confused and this problem is due in 10 min :/

OpenStudy (mathlegend):

3x(5x-1) Then you would factor out -10 and 2

OpenStudy (mathlegend):

-2(5x-1)

OpenStudy (mathlegend):

So combining those you would get 3x(5x-1)-2(5x-1)

OpenStudy (mathlegend):

Now you take the numbers outside of the parentheses and put those together and the inside comes after. (3x-2)(5x-1)

OpenStudy (mathlegend):

That is your answer....

OpenStudy (anonymous):

is says thast wrong when i type it in

OpenStudy (mathlegend):

What are you typing it into?

OpenStudy (anonymous):

xyz.com im doing homework

OpenStudy (anonymous):

xyz homework

OpenStudy (mathlegend):

Okay what do they say is the answer?

OpenStudy (anonymous):

they dont tell you it just says wrong

OpenStudy (mathlegend):

Okay one sec

OpenStudy (anonymous):

kk

OpenStudy (mathlegend):

Okay common mistake

OpenStudy (mathlegend):

I'll give you the answer to submit

OpenStudy (mathlegend):

( 5x - 1 )( 3x - 2 )

OpenStudy (anonymous):

thanks you for all the help, sorry i didnt quite catch on to this problem

OpenStudy (mathlegend):

Does it say that is right?

OpenStudy (anonymous):

yea

OpenStudy (mathlegend):

You know the strange thing is I gave you (3x-2)(5x-1) which is actually the same as (5x-1)(3x-2) That computer must not know about the Commutative Property

OpenStudy (mathlegend):

Glad I could help though!